OpenStudy (anonymous):

soooo lost Solve 4x + 3y = 13 by the substitution method. x + y = 4

OpenStudy (smurfy14):

y=3 x=1

OpenStudy (anonymous):

For substitution, you need to solve one equation in terms of the other. The easiest one would be the second one. So subtract over the x giving: y=4-x. From here you SUBSTITUTE it into the other one. 4x+3(4-x)=13 Distribute: 4x+12-3x=13 combine. x+12=13. Subtract x=1. From there, take that and substitute that into what you solved y. y=4-1=3. :)

OpenStudy (anonymous):

\[y=3 \ or\ y=1\]

OpenStudy (anonymous):

Hopefully my explanation helps instead of just giving out answers.

OpenStudy (smurfy14):

solve for x: x=-y+4 then plug that in to other one: 4(-y+4)+3y=13

OpenStudy (anonymous):

x= 4-y frm second eqn putting in first 4(4-y) + 3y = 13 Solve this to get value of y put in x+y = 4 and get value of x

OpenStudy (smurfy14):

then once you find y go back and plug it in to find x. x+3=4 so x=1